# The algebraic and geometric classification of

### v1

@@ 1.2. $2$-dimensional pre-Lie algebras
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{01})$, then $\phi(e_{1})=e_{1}+\nu e_{2}$ and $\phi(e_{2})=e_{2}$, for $\nu\in\mathbb{C}$.
@@
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{02})$, then $\phi(e_{1})=e_{1}+\nu e_{2}$ and $\phi(e_{2})=e_{2}$, for $\nu\in\mathbb{C}$.
@@
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{03})$, then $\phi(e_{1})=\xi e_{1}+\nu e_{2}$ and $\phi(e_{2})=\xi^{2}e_{2}$, for $\xi\in\mathbb{C}^{\ast}$ and $\nu\in\mathbb{C}$.
@@
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{04})$, then $\phi(e_{1})=\xi e_{1}$ and $\phi(e_{2})=e_{2}$, for $\xi\in\mathbb{C}^{\ast}$.
@@
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{05}^{\alpha\neq 1})$, then $\phi(e_{1})=e_{1}$ and $\phi(e_{2})=\xi e_{2}$, for $\xi\in\mathbb{C}^{\ast}$.
@@
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{05}^{1})$, then $\phi(e_{1})=e_{1}+\nu e_{2}$ and $\phi(e_{2})=\xi e_{2}$, for $\xi\in\mathbb{C}^{\ast}$ and $\nu\in\mathbb{C}$.
@@
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{06}^{\alpha\neq 0})$, then $\phi(e_{1})=e_{1}$ and $\phi(e_{2})=\xi e_{2}$, for $\xi\in\mathbb{C}^{\ast}$.
@@
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{06}^{0})$, then $\phi(e_{1})=e_{1}+\nu e_{2}$ and $\phi(e_{2})=\xi e_{2}$, for $\xi\in\mathbb{C}^{\ast}$ and $\nu\in\mathbb{C}$.
@@
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{07})$, then $\phi\in\mathbb{S}_{2},$ i.e. $\phi(e_{1})=e_{1},\ \phi(e_{2})=e_{2}$ or $\phi(e_{1})=e_{2},\ \phi(e_{2})=e_{1}.$
@@
1.3. The algebraic classification of compatible pre-Lie algebras
@@
$\mathcal{C}_{01}$ :::: $e_{1}\ast e_{1}=e_{1}+e_{2},\ e_{2}\ast e_{1}=e_{2}.$
$\mathcal{C}_{02}$ :::: $e_{1}\ast e_{1}=e_{1}+e_{2},\ e_{1}\ast e_{2}=e_{2}.$
$\mathcal{C}_{03}$ :::: $e_{1}\ast e_{1}=e_{2}.$
$\mathcal{C}_{04}$ :::: $e_{2}\ast e_{1}=e_{1}.$
$\mathcal{C}_{05}^{\alpha}$ :::: $e_{1}\ast e_{1}=e_{1},\ e_{1}\ast e_{2}=\alpha e_{2}.$
$\mathcal{C}_{06}^{\alpha}$ :::: $e_{1}\ast e_{1}=e_{1},\ e_{1}\ast e_{2}=\alpha e_{2},\ e_{2}\ast e_{1}=e_{2}.$
$\mathcal{C}_{07}$ :::: $e_{1}\ast e_{1}=e_{1},\ e_{2}\ast e_{2}=e_{2}.$
$\mathcal{C}_{08}$ :::: $e_{1}\ast e_{1}=e_{1},\ e_{1}\ast e_{2}=2e_{2},\ e_{2}\ast e_{1}=\frac{1}{2}e_{1}+e_{2},\ e_{2}\ast e_{2}=e_{2}.$
$\mathcal{C}_{09}^{\alpha,\beta\neq 0}$ :::: {$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\beta e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2}.$\\ \end{tabular}\right.$
$\mathcal{C}_{10}^{\alpha,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{11}^{\alpha,\beta\neq\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{1}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{12}^{\alpha,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{1}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{1}\ast e_{2}=\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{13}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{14}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=e_{1}+\alpha e_{2},$&$e_{2}\ast e_{1}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{15}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{16}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=e_{1},$&$e_{1}\ast e_{2}=\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{17}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=e_{1}+\alpha e_{2},$&$e_{1}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{18}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&&&\\ $e_{1}\ast e_{1}=\alpha e_{2},$&$e_{1}\ast e_{2}=e_{1},$&$e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{19}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&&\\ $e_{1}\ast e_{1}=\alpha e_{2},$&$e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=2e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{20}^{\alpha,\beta}$ :::: {$e_{2}\cdot e_{1}=e_{1},$&\\ $e_{2}\ast e_{1}=\alpha e_{1},$&$e_{2}\ast e_{2}=e_{1}+\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{21}^{\alpha,\beta}$ :::: {$e_{2}\cdot e_{1}=e_{1},$&\\ $e_{2}\ast e_{1}=\alpha e_{1},$&$e_{2}\ast e_{2}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{22}^{\alpha\neq 0,\beta}$ :::: {$e_{2}\cdot e_{1}=e_{1},$&&\\ $e_{1}\ast e_{2}=\alpha e_{1},$&$e_{2}\ast e_{1}=\beta e_{1},$&$e_{2}\ast e_{2}=e_{1}+\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{23}^{\alpha\neq 0,\beta}$ :::: {$e_{2}\cdot e_{1}=e_{1},$&&\\ $e_{1}\ast e_{2}=\alpha e_{1},$&$e_{2}\ast e_{1}=\beta e_{1},$&$e_{2}\ast e_{2}=\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{24}^{\alpha,\beta,\gamma}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\alpha e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1}+e_{2},$&$e_{1}\ast e_{2}=\gamma e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{25}^{\alpha,\beta,\gamma}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\alpha e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1},$&$e_{1}\ast e_{2}=\gamma e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{26}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\frac{1}{2}e_{2},$&\\ $e_{1}\ast e_{1}=2\alpha e_{1},$&$e_{1}\ast e_{2}=\alpha e_{2},$&$e_{2}\ast e_{2}=e_{1}.$\end{tabular}\right.$
$\mathcal{C}_{27}^{\alpha,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\frac{1}{2}e_{2},$&&\\ $e_{1}\ast e_{1}=2\alpha e_{1},$&$e_{1}\ast e_{2}=e_{1}+\alpha e_{2},$&$e_{2}\ast e_{1}=2e_{1}$&$e_{2}\ast e_{2}=\beta e_{1}+e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{28}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$\\ $e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{29}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{30}^{\alpha,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&&&\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{1}\ast e_{2}=e_{1},$&$e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{31}^{\alpha,\beta,\gamma}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\alpha e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1}+e_{2},$&$e_{1}\ast e_{2}=\gamma e_{2},$&$e_{2}\ast e_{1}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{32}^{\alpha,\beta,\gamma}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\alpha e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1},$&$e_{1}\ast e_{2}=\gamma e_{2},$&$e_{2}\ast e_{1}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{33}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{2}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{34}^{\alpha,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\alpha e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2},$&$e_{2}\ast e_{2}=e_{1}+\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{35}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\alpha e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{36}^{\alpha,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{1}\ast e_{2}=2\alpha e_{2},$&$e_{2}\ast e_{1}=e_{1}+\alpha e_{2},$&$e_{2}\ast e_{2}=2e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{37}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=e_{1}+2\alpha e_{2},$&$e_{2}\ast e_{1}=2e_{1}+\alpha e_{2},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{38}^{\alpha,\beta,\gamma}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{2}=e_{2},$&&\\ $e_{1}\ast e_{1}=\left(\gamma+\beta-\alpha\right)e_{1}-\beta e_{2},$&$e_{1}\ast e_{2}=\alpha e_{1}+\beta e_{2},$\\ $e_{2}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{2}\ast e_{2}=-\alpha e_{1}+\gamma e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{39}^{\alpha,\beta\neq\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{2}\ast e_{2}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{40}^{\alpha,\beta,\gamma}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$&$e_{2}\cdot e_{1}=\frac{1}{2}e_{1}+e_{2},$&$e_{2}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{1}\ast e_{2}=2\alpha e_{2},$&$e_{2}\ast e_{1}=\gamma e_{1}+\alpha e_{2},$&$e_{2}\ast e_{2}=2\gamma e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{41}^{\alpha,\beta\neq 0}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$&\\ $e_{2}\cdot e_{1}=\frac{1}{2}e_{1}+e_{2},$&$e_{2}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\beta e_{1}+2\alpha e_{2},$&\\ $e_{2}\ast e_{1}=\left(2\beta+\frac{1}{2}\alpha\right)e_{1}+\alpha e_{2},$&$e_{2}\ast e_{2}=\frac{1}{2}\beta e_{1}+\left(\alpha+\beta\right)e_{2}.$\end{tabular}\right.$
@@
$\mathcal{C}_{24}^{1,\beta,\gamma}$cong $\mathcal{C}_{25}^{1,\beta,\gamma},$ $\mathcal{C}_{31}^{0,\beta,\gamma}$cong $\mathcal{C}_{32}^{0,\beta,\gamma},$ $\mathcal{C}_{38}^{\alpha,\beta,\gamma}$cong $\mathcal{C}_{38}^{\beta,\alpha,-\alpha+\beta+\gamma},$
$\mathcal{C}_{39}^{\alpha,\beta}$cong $\mathcal{C}_{39}^{\beta,\alpha},$ $\mathcal{C}_{40}^{\alpha,\beta,\gamma}$cong $\mathcal{C}_{40}^{4\gamma-\alpha,\beta-8\alpha+16\gamma,\gamma},$ $\mathcal{C}_{41}^{\alpha,\beta}$cong $\mathcal{C}_{41}^{\alpha+4\beta,-\beta}.$
1.3.1. Compatible pre-Lie algebras defined on $\mathcal{C}_{01}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{01},\mathcal{C}_{01})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{01}$ are of the form:
{$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{3}e_{2},$&$e_{2}\ast e_{1}=\alpha_{1}e_{2}.$\end{tabular}\right.\qquad$
Then we have the following cases:
  • If $\alpha_{3}\neq 0$, then choose $\nu=-\frac{\alpha_{2}}{\alpha_{3}}$ and obtain the parametric family:
    $\mathcal{C}_{09}^{\alpha,\beta\neq 0}:=\left\{\begin{tabular}[]{lll}$e_{1}\cdot e_{1}=e_{1}+e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\beta e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2}.$\end{tabular}\right.\qquad$

### v2

@@ 1. $(1)$
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{01})$, then $\phi(e_{1})=e_{1}+\nu e_{2}$ and $\phi(e_{2})=e_{2}$ for $\nu\in\mathbb{C}$.
@@ 2. $(2)$
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{02})$, then $\phi(e_{1})=e_{1}+\nu e_{2}$ and $\phi(e_{2})=e_{2}$ for $\nu\in\mathbb{C}$.
@@ 3. $(3)$
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{03})$, then $\phi(e_{1})=\xi e_{1}+\nu e_{2}$ and $\phi(e_{2})=\xi^{2}e_{2}$ for $\xi\in\mathbb{C}^{\ast}$ and $\nu\in\mathbb{C}$.
@@ 4. $(4)$
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{04})$, then $\phi(e_{1})=\xi e_{1}$ and $\phi(e_{2})=e_{2}$ for $\xi\in\mathbb{C}^{\ast}$.
@@ 5. $(5)$
    If $\phi\in\mathrm{Aut}\bigl{(}\mathcal{C}_{05}^{\alpha\neq 1}\bigr{)}$, then $\phi(e_{1})=e_{1}$ and $\phi(e_{2})=\xi e_{2}$ for $\xi\in\mathbb{C}^{\ast}$.
@@ 6. $(6)$
    If $\phi\in\mathrm{Aut}\bigl{(}\mathcal{C}_{05}^{1}\bigr{)}$, then $\phi(e_{1})=e_{1}+\nu e_{2}$ and $\phi(e_{2})=\xi e_{2}$ for $\xi\in\mathbb{C}^{\ast}$ and $\nu\in\mathbb{C}$.
@@ 7. $(7)$
    If $\phi\in\mathrm{Aut}\bigl{(}\mathcal{C}_{06}^{\alpha\neq 0}\bigr{)}$, then $\phi(e_{1})=e_{1}$ and $\phi(e_{2})=\xi e_{2}$ for $\xi\in\mathbb{C}^{\ast}$.
@@ 8. $(8)$
    If $\phi\in\mathrm{Aut}\bigl{(}\mathcal{C}_{06}^{0}\bigr{)}$, then $\phi(e_{1})=e_{1}+\nu e_{2}$ and $\phi(e_{2})=\xi e_{2}$ for $\xi\in\mathbb{C}^{\ast}$ and $\nu\in\mathbb{C}$.
@@ 9. $(9)$
    If $\phi\in\mathrm{Aut}(\mathcal{C}_{07})$, then $\phi\in\mathbb{S}_{2}$, i.e., $\phi(e_{1})=e_{1}$, $\phi(e_{2})=e_{2}$ or $\phi(e_{1})=e_{2}$, $\phi(e_{2})=e_{1}.$
@@ 10. $(10)$
2.3 The algebraic classification of compatible pre-Lie algebras
@@
$\mathcal{C}_{01}\colon\ e_{1}\ast e_{1}=e_{1}+e_{2},\quad e_{2}\ast e_{1}=e_{2},$
$\mathcal{C}_{02}\colon\ e_{1}\ast e_{1}=e_{1}+e_{2},\quad e_{1}\ast e_{2}=e_{2},$
$\mathcal{C}_{03}\colon\ e_{1}\ast e_{1}=e_{2},$
$\mathcal{C}_{04}\colon\ e_{2}\ast e_{1}=e_{1},$
$\mathcal{C}_{05}^{\alpha}\colon\ e_{1}\ast e_{1}=e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2},$
$\mathcal{C}_{06}^{\alpha}\colon\ e_{1}\ast e_{1}=e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2},\quad e_{2}\ast e_{1}=e_{2},$
$\mathcal{C}_{07}\colon\ e_{1}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{08}\colon\ e_{1}\ast e_{1}=e_{1},\quad e_{1}\ast e_{2}=2e_{2},\quad e_{2}\ast e_{1}=\tfrac{1}{2}e_{1}+e_{2},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{13}^{1}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=e_{1},\quad e_{1}\ast e_{2}=e_{2},\quad e_{2}\ast e_{1}=e_{2},$
$\mathcal{C}_{15}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{2},$
$\mathcal{C}_{16}^{0}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=e_{1},$
$\mathcal{C}_{18}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{2},\quad e_{1}\ast e_{2}=e_{1},\quad e_{2}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{24}^{0,\beta,0}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{1}=\beta e_{1}+e_{2},$
$\mathcal{C}_{25}^{0,\beta,0}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{1}=\beta e_{1},$
$\mathcal{C}_{29}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{1}=\alpha e_{1},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{30}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},\quad e_{1}\ast e_{2}=e_{1},\quad e_{2}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{31}^{1,\beta,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\beta e_{1}+e_{2},\quad e_{1}\ast e_{2}=\beta e_{2},\quad e_{2}\ast e_{1}=\beta e_{2},$
$\mathcal{C}_{32}^{1,\beta,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\beta e_{1},\quad e_{1}\ast e_{2}=\beta e_{2},\quad e_{2}\ast e_{1}=\beta e_{2},$
$\mathcal{C}_{34}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2},\quad e_{2}\ast e_{1}=\alpha e_{2},\quad e_{2}\ast e_{2}=e_{1}+\beta e_{2},$
$\mathcal{C}_{35}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2},\quad e_{2}\ast e_{1}=\alpha e_{2},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{38}^{\alpha,\beta,\gamma}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{2}=e_{2},\quad e_{1}\ast e_{1}=(\gamma+\beta-\alpha)e_{1}-\beta e_{2},$
$\ e_{1}\ast e_{2}=\alpha e_{1}+\beta e_{2},\quad e_{2}\ast e_{1}=\alpha e_{1}+\beta e_{2},\quad e_{2}\ast e_{2}=-\alpha e_{1}+\gamma e_{2},$
$\mathcal{C}_{39}^{\alpha,\beta\neq\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{2}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1},\quad e_{2}\ast e_{2}=\beta e_{2}.$
@@
$\mathcal{C}_{24}^{1,\beta,\gamma}\cong\mathcal{C}_{25}^{1,\beta,\gamma},\quad\mathcal{C}_{31}^{0,\beta,\gamma}\cong\mathcal{C}_{32}^{0,\beta,\gamma},\quad\mathcal{C}_{38}^{\alpha,\beta,\gamma}\cong\mathcal{C}_{38}^{\beta,\alpha,-\alpha+\beta+\gamma},$
$\mathcal{C}_{39}^{\alpha,\beta}\cong\mathcal{C}_{39}^{\beta,\alpha},\quad\mathcal{C}_{40}^{\alpha,\beta,\gamma}\cong\mathcal{C}_{40}^{4\gamma-\alpha,\beta-8\alpha+16\gamma,\gamma},\quad\mathcal{C}_{41}^{\alpha,\beta}\cong\mathcal{C}_{41}^{\alpha+4\beta,-\beta}.$
2.3.1 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{01}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{01},\mathcal{C}_{01})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{01}$ are of the form $e_{1}\cdot e_{1}=e_{1}+e_{2}$, $e_{2}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$, $e_{2}\ast e_{1}=\alpha_{1}e_{2}$. Then we have the following cases:
  • If $\alpha_{3}\neq 0$, then choose $\nu=-\frac{\alpha_{2}}{\alpha_{3}}$ and obtain the parametric family
    $\mathcal{C}_{09}^{\alpha,\beta\neq 0}\colon\ e_{1}\cdot e_{1}=e_{1}+e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
[…]